OpenStudy (anonymous):

A=(1/2)b(h_1+h_2) solve for h_2

OpenStudy (jim766):

A = 1/2 b(h1 + h2) lets start by getting rid of the 1/2, mult both sides by 2

OpenStudy (jim766):

2A = b(h1 + h2) now how can we get the b out of there?

OpenStudy (anonymous):

\[\frac{ 2A }{ b }-h _{1}=h_{2}\]

OpenStudy (anonymous):

(2A-bh_1)/b=h_2

OpenStudy (anonymous):

i think @v4xN0s is wrong

OpenStudy (jim766):

nope he is right... 2A = b(h1 + h2) div both sides by b 2a/b = h1 + h2

OpenStudy (jim766):

then subtract h1 from both sides 2a/b - h1 = h2

OpenStudy (anonymous):

so I'm wrong?

OpenStudy (phi):

(2A-bh_1)/b=h_2 is the same as 2A/b -bh_1/b = h_2 or 2A/b - h_1 = h_2

OpenStudy (anonymous):

nevermind we both right

OpenStudy (phi):

It is good to notice they are the same, in case a multiple choice does not have your version, and you can recognize the same answer in a different form